# thanks: Corresponding author

### v1

@@ III.2 Isothermal Process
The isothermal process corresponds to the $P$-$V$ ensemble commonly employed in black hole thermodynamics. Here the only variable parameters are $M$ and $P$, which are related through the isothermal constraint $T(r_{h};M,P)=T_{0}$. Therefore, we take $\lambda=(M,P;\lambda_{\mathrm{fixed}})$. Substituting this into the criterion (21) and choosing $r_{h}$ as the path parameter gives (see Appendix C for justification)
@@
In general, this relation has no simple closed form. However, for photons ($\epsilon=0$), many black-hole metrics satisfy $\partial_{P}\Phi_{0}=0$, for example when the pressure enters the metric only through a term proportional to $Pr^{2}$. In such cases, the criterion (38) reduces to
@@
This result is consistent with the criterion for the isobaric process, but differs in that under the isobaric constraint the first law of thermodynamics requires ${\mathrm{d}M}/{\mathrm{d}r_{h}}\neq 0$, ensuring that $r_{\mathrm{ps}}(r_{h})$ remains monotonic. In contrast, under the isothermal constraint the first law takes the form
@@
indicating that ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ may vanish, which in turn allows $r_{\mathrm{ps}}(r_{h})$ to become non-monotonic. This behavior is physically intuitive: when ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$, one must have
@@
This means that, near an extremum of $M(r_{h})$, the same $M$ corresponds to two distinct $P$. As $r_{\rm ps}$ is $P$-independent (since $\partial_{P}\Phi_{0}=0$) while $r_{h}$ depends on $P$, consequently a single $r_{\rm ps}$ maps to two different $r_{h}$, and $r_{\rm ps}(r_{h})$ is necessarily non-monotonic. Next, we apply the criterion (38) to the metric (23).
@@ III.2.1 Photon sphere ($\epsilon=0$)
$\lim_{r_{h}\to 0^{+}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=-\infty,\qquad\lim_{r_{h}\to\infty}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=+\infty,$ (46)
hence there exists a unique $r_{h0}$ such that $\bigl.{\mathrm{d}M}/{\mathrm{d}r_{h}}\bigr|_{r_{h0}}=0$. By the criterion (39), it follows that:
@@
  • $q\neq 0$: $r_{\rm ps}(r_{h})$ is non-monotonic. It decreases for small $r_{h}$, reaches a single minimum at $r_{h0}$, and then increases.
This result reveals that the presence of electric charge $q$ breaks the monotonic behavior of $r_{\mathrm{ps}}(r_{h})$, making it globally non-monotonic–an aspect that has been omitted in previous studies. The explicit expression for ${dr_{\mathrm{ps}}}/{dr_{h}}$ reads
@@
For $d=4$ (the RN-AdS case), this reduces to
@@
and, in the Sch-AdS limit ($q=0$), further simplifies to
@@
This coincides with the standard Sch-AdS result, from which ${\mathrm{d}r_{\rm ps}}/{\mathrm{d}r_{h}}>0$ is immediately evident.
Figure 4 presents the numerically computed profiles of $r_{\mathrm{ps}}(r_{h})$. For the Sch-AdS black hole, $r_{\mathrm{ps}}(r_{h})$ begins at $r_{h,\min}=(d-3)/(4\pi T_{0})$ and increases monotonically thereafter, remaining globally monotonic in full agreement with Eq. (49). In contrast, the RN-AdS black hole exhibits the expected non-monotonic behavior: although ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$ at the extremum, the surface gravity remains positive ($\kappa=2\pi T_{0}>0$), indicating that the black hole is nonextremal and the curve need not be truncated.
The breakdown of monotonicity naturally raises an important question: does it affect the discontinuous jump of $r_{\mathrm{ps}}$ across the phase transition, potentially obscuring its optical signature?
The answer is: possibly. If the small and large black hole radii, $r_{h,s}$ and $r_{h,l}$, lie on opposite sides of the extremum of $r_{\mathrm{ps}}(r_{h})$, the amplitude of the discontinuous change in $r_{\mathrm{ps}}$ may be reduced, requiring higher observational resolution to detect it. In the extreme case where $r_{\mathrm{ps}}(r_{h,s})=r_{\mathrm{ps}}(r_{h,l})$, the discontinuity would vanish entirely, rendering the phase transition optically unobservable.
From a thermodynamic perspective, however, this is not necessarily a drawback. As shown earlier, $r_{\mathrm{c}}(r_{h})$ remains monotonic in the isobaric process, while $r_{\mathrm{ps}}(r_{h})$ becomes non-monotonic in the isothermal one. Such a contrast in optical behavior between the two ensembles offers a potential diagnostic for identifying the underlying thermodynamic process of a black hole– a perspective that provides a new avenue for probing black hole thermodynamics, to be explored in future work.
Another possibility is that $r_{h,s}$ and $r_{h,l}$ lie on the same side of the extremum of $r_{\mathrm{ps}}(r_{h})$, so that $r_{\mathrm{ps}}(r_{h})$ remains locally monotonic throughout the phase-transition region. This scenario indeed occurs for the RN-AdS black hole studied here, and in Appendix F we provide an explicit proof for the $d=4$ case.
Figure 5 shows the numerically computed profiles of $P(r_{h})$ and $P(r_{\mathrm{ps}})$ at fixed parameters, where the characteristic S-shaped behavior of the phase transition is fully preserved. In this configuration, one can observe both the global non-monotonicity of $r_{\mathrm{ps}}(r_{h})$ and the discontinuous jump of $r_{\mathrm{ps}}$ across the transition. The coexistence of these two optical signatures provides an especially powerful means for extracting thermodynamic information from black holes.
Figure 4: Behavior of $r_{\rm ps}(r_{h})$ along isothermal paths in $d=4$. The marked point on the curve corresponds to $P=0$ at $r_{h_{\min}}=(d-3)/(4\pi T_{0})$, and the unphysical portion with $P<0$ is truncated. For the Sch-AdS case, the temperature is likewise set to the value $T=0.8\,T_{\mathrm{c}}$ corresponding to the $q=1$ reference point. The same choice is adopted for subsequent figures and will not be repeated.
Figure 5: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm ps})$ for $d=4$ and $q=1$.
@@ III.2.2 ISCO ($\epsilon=1$)
where $\partial_{P}\Phi_{1}>0$ (see Appendix D for proof). Since van der Waals-type behavior occurs only for charged black holes ($q\neq 0$), we analyze the RN-AdS and Sch-AdS ($q=0$) cases separately. Moreover, because the explicit expression for ${\mathrm{d}r_{\mathrm{c}}}/{\mathrm{d}r_{h}}$ is algebraically cumbersome, we focus on whether it changes sign rather than displaying the full formula.
(i) Sch-AdS ($q=0$)
@@
It has a single zero at $r_{h1}={(d-3)}/{2\pi T_{0}}$, with ${\mathrm{d}P}/{\mathrm{d}r_{h}}>0$ for $r_{h}<r_{h1}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$ for $r_{h}>r_{h1}$. Using the criterion (38), together with ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}<0$, $\Gamma_{1}>0$, and $\partial_{P}\Phi_{1}>0$, one thus expects $r_{\rm isco}(r_{h})$ to first decrease and then increase. To make this more explicit, we rewrite the criterion (38) as
@@
For sufficiently large $r_{h}$ one has ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$, which forces $\delta<0$. It remains to verify that $\delta>0$ for small $r_{h}$, ensuring a single zero crossing. In this spirit we find
$\begin{cases}\delta\bigl(3r_{h_{\min}},\,r_{h_{\min}},\,4,\,0,\,T_{0}\bigr)=\dfrac{1232\,\pi^{2}T_{0}^{2}}{27}>0,&d=4,\\[6.0pt] \delta\bigl(+\infty,\,r_{h_{\min}},\,d,\,0,\,T_{0}\bigr)=\dfrac{512\,\pi^{3}T_{0}^{3}\,r_{\rm isco}}{(d-3)^{2}(d-1)}>0,&d>4,\end{cases}$ (54)
where $r_{h_{\min}}=(d-3)/(4\pi T_{0})$ is defined by $P(r_{h_{\min}})=0$. For $d=4$ and $P=0$, the metric (23) reduces to the Schwarzschild case, giving $r_{\rm isco}=3\,r_{h_{\min}}$. For $d>4$ and $P\to 0$, the higher-dimensional Schwarzschild black hole has no ISCO, hence $r_{\rm isco}\to+\infty$. Equation (54) then confirms that $\delta$ indeed crosses zero. Consequently, under the isothermal constraint, the function $r_{\rm isco}(r_{h})$ is non-monotonic.
(ii) RN-AdS ($q\neq 0$)
Since the RN-AdS black hole exhibits a van der Waals–like phase transition, one has
@@
where the equality holds precisely at the critical point $(T_{c},P_{c},r_{hc})$. According to the criterion (38), when $r_{h}\geq r_{h0}$ (where $r_{h0}$ is the zero of $\mathrm{d}M/\mathrm{d}r_{h}$), it follows that $\delta<0$ and hence $\mathrm{d}r_{\mathrm{isco}}/\mathrm{d}r_{h}>0$, so $r_{\mathrm{isco}}(r_{h})$ increases monotonically. For $r_{h}<r_{h0}$, however, $\delta$ may become positive. To determine this behavior, we examine the limit $r_{h}\to 0^{+}$.
@@
Meanwhile, from Appendix E, as $r_{h}\to 0^{+}$ one has
@@
This demonstrates that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic over the full domain of $r_{h}$.
The same conclusion holds for $T<T_{c}$. Along both the small–black–hole and large–black–hole stable branches, one still has $\mathrm{d}P/\mathrm{d}r_{h}<0$, and the leading-order behavior as $r_{h}\to 0^{+}$ remains identical to the $T\geq T_{c}$ case above. The only difference is that for $T<T_{c}$ an intermediate spinodal (unstable) branch appears, on which $\mathrm{d}P/\mathrm{d}r_{h}>0$. Across this segment, the sign of $\delta$ may reverse, thereby altering the local monotonicity of $r_{\mathrm{isco}}(r_{h})$. In all cases, however, under the isothermal constraint $r_{\mathrm{isco}}(r_{h})$ is globally non-monotonic.
Figure 6 presents the numerically computed profiles of $r_{\mathrm{isco}}(r_{h})$. It can be seen that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic for both charged and uncharged cases. The distinction lies in that, for the uncharged (Sch-AdS) case, $r_{\mathrm{isco}}(r_{h})$ starts at $r_{h_{\min}}=(d-3)/(4\pi T_{0})$ where $P(r_{h})=0$, whereas for the charged (RN-AdS) black hole, $P(r_{h})>0$ holds throughout the entire domain.
Appendix F likewise shows that $r_{h,s}$ and $r_{h,l}$ lie on the same side of the extremum of $r_{\mathrm{isco}}(r_{h})$. Consequently, $P(r_{\mathrm{isco}})$ preserves the full local SS–shaped structure of $P(r_{h})$ across the phase transition. An explicit example for fixed parameters is shown in Fig. 7.
@@
Figure 7: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm isco})$ for $d=4$ and $q=1$.
It is worth noting that, according to the first law of thermodynamics, a non-monotonic $M(r_{h})$ implies that the same mass $M$ can correspond to two distinct pressures $P$. Since both $r_{\mathrm{isco}}$ and $r_{h}$ depend on $P$, one might expect $r_{\mathrm{isco}}(r_{h})$ to be less prone to non-monotonic behavior. However, our results show that non-monotonicity indeed occurs, even for the Sch-AdS black hole.
To better understand this, we examine the criterion (38). Because $\partial_{M}\Phi_{1}<0$ and $\partial_{P}\Phi_{1}>0$, $r_{\mathrm{isco}}(r_{h})$ is guaranteed to be monotonic as long as ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}$ have opposite signs. Yet the first law,
requires that ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}$ share the same sign in certain ranges of $r_{h}$.
For the $d$-dimensional RN-AdS black hole, ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ is non-monotonic and can take negative values. Since the first law demands ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$, the two derivatives become of the same sign, allowing the criterion (38) to cross zero. For the $d$-dimensional Sch-AdS black hole, comparing
$T_{0}S^{\prime}(r_{h})=\frac{4\pi}{c_{d}}\,(d-2)\,T_{0}\,r_{h}^{\,d-3},$ (61)
with
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=\frac{2(d-3)\,r_{h}^{\,d-4}}{(d-1)c_{d}}+\frac{4\pi(d-2)T_{0}\,r_{h}^{\,d-3}}{(d-1)c_{d}},$ (62)
immediately reveals their relative size in the small-horizon regime. For instance, at the minimal value $r_{h_{\min}}=(d-3)/(4\pi T_{0})$, one finds
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}$ $=\frac{d}{d-1}\frac{(4\pi)^{4-d}T_{0}}{c_{d}}(\frac{d-3}{T_{0}})^{d-3},$ (63)
$T_{0}S^{\prime}(r_{h})$ $=(d-2)\frac{(4\pi)^{4-d}T_{0}}{c_{d}}(\frac{d-3}{T_{0}})^{d-3}.$ (64)
It is clear that
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}<T_{0}S^{\prime}(r_{h}),$
which, through the first law $\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{0}S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ requires $\frac{\mathrm{d}P}{\mathrm{d}r_{h}}>0$. Thus ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ acquire the same sign, allowing the criterion (38) to cross zero.

### v2

@@ III.2 Isothermal Process
The isothermal process corresponds to the $P$-$V$ ensemble commonly employed in black hole thermodynamics. Here, the only variable parameters are $M$ and $P$, which are related through the isothermal constraint $T(r_{h};M,P)=T_{0}$. Therefore, we take $\lambda=(M,P;\lambda_{\mathrm{fixed}})$. Substituting this into the criterion (21) and choosing $r_{h}$ as the path parameter yields the following result [see Appendix D for justification of the existences of $M(r_{h})$ and $P(r_{h})$]
@@
In general, this relation has no simple closed form. However, for photons ($\epsilon=0$), many black hole metrics satisfy $\partial_{P}\Phi_{0}=0$, e.g., when the pressure enters the metric only through a term proportional to $Pr^{2}$. In such cases, the criterion (38) reduces to
@@
This result is consistent with the criterion for the isobaric process, but differs in that the isobaric constraint requires ${\mathrm{d}M}/{\mathrm{d}r_{h}}>0$ via the first law, ensuring the monotonicity of $r_{\mathrm{ps}}(r_{h})$. In contrast, under the isothermal constraint, the first law takes the form
@@
indicating that ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ may vanish or even become negative, which in turn allows $r_{\mathrm{ps}}(r_{h})$ to become nonmonotonic. This behavior is physically intuitive: when ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$, one must have
@@
This means that, near an extremum of $M(r_{h})$, the same $M$ corresponds to two distinct $P$. Since $r_{\rm ps}$ is $P$-independent (since $\partial_{P}\Phi_{0}=0$) while $r_{h}$ depends on $P$, a single $r_{\rm ps}$ maps to two different $r_{h}$, and thus $r_{\rm ps}(r_{h})$ is necessarily nonmonotonic.
For massive particles ($\epsilon=1$), one typically has $\partial_{P}\Phi_{1}\neq 0$. Taking into account that $M$ and $P$ usually affect $r_{\rm isco}$ in opposite ways, we assume that $\partial_{M}\Phi_{1}$ and $\partial_{P}\Phi_{1}$ carry opposite signs (as is indeed the case for RN–AdS black holes). Under this assumption, the criterion (38) shows that if $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ have opposite signs, then $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ cannot cross zero, so $r_{\rm isco}(r_{h})$ remains monotonic. Conversely, if $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ share the same sign, $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$ can in principle vanish, leading to a nonmonotonic $r_{\rm isco}(r_{h})$. Interestingly, the first law naturally allows for such nonmonotonic behavior: from Eq. (41) one sees that $\mathrm{d}P/\mathrm{d}r_{h}>0$ implies $\mathrm{d}M/\mathrm{d}r_{h}>0$, while $\mathrm{d}M/\mathrm{d}r_{h}<0$ implies $\mathrm{d}P/\mathrm{d}r_{h}<0$, so there exist regimes where $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ have the same sign and $r_{\rm isco}(r_{h})$ can become nonmonotonic.
Conversely, if we assume that $\partial_{M}\Phi_{1}$ and $\partial_{P}\Phi_{1}$ share the same sign, the first law will require that $r_{\rm isco}(r_{h})$ be nonmonotonic. The reason is simple: At an extremum of $M(r_{h})$, one has $\mathrm{d}M/\mathrm{d}r_{h}=0$, and thus Eq. (41) requires $\mathrm{d}P/\mathrm{d}r_{h}<0$. Substituting these relations into the criterion (38) yields
$\left.\mathrm{sgn}\!\left(\frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}\right)\right|_{\mathrm{d}M/\mathrm{d}r_{h}=0}=\mathrm{sgn}\!\left(-\frac{1}{\Gamma_{1}}\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{1}\right).$ (43)
At an extremum of $P(r_{h})$, one instead has $\mathrm{d}P/\mathrm{d}r_{h}=0$, and thus Eq. (41) gives $\mathrm{d}M/\mathrm{d}r_{h}>0$. Inserting this into the criterion (38) leads to
$\left.\mathrm{sgn}\!\left(\frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}\right)\right|_{\mathrm{d}P/\mathrm{d}r_{h}=0}=\mathrm{sgn}\!\left(-\frac{1}{\Gamma_{1}}\,\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}\right).$ (44)
It is easy to see that the two above expressions have different signatures. Therefore, the first law enforces a sign change in $\mathrm{d}r_{\rm isco}/\mathrm{d}r_{h}$, implying that $r_{\rm isco}(r_{h})$ is necessarily nonmonotonic, with its extremum located between those of $M(r_{h})$ and $P(r_{h})$. Whether such a configuration is actually realized depends on the specific black hole model and must be examined case by case.
In any case, these analyses show that, in the isothermal ensemble, the first law no longer enforces $r_{c}(r_{h})$ to be monotonic, but instead allows nonmonotonic behavior. Once such nonmonotonicity arises, the jump $\Delta_{c}=r_{c}(r_{h,l})-r_{c}(r_{h,s})$ can be strongly suppressed or even vanish, thereby undermining the reliability of circular orbit radii as encoders of phase transitions. Next, we apply criterion (38) to the metric (23) to verify the above analysis explicitly.
@@ III.2.1 Photon sphere ($\epsilon=0$)
$\lim_{r_{h}\to 0^{+}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=-\infty,\qquad\lim_{r_{h}\to+\infty}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=+\infty,$ (48)
hence there exists a unique $r_{h,e_{1}}$ such that $\bigl.{\mathrm{d}M}/{\mathrm{d}r_{h}}\bigr|_{r_{h,e_{1}}}=0$. By the criterion (39), it follows that:
@@
  • $q\neq 0$: $r_{\rm ps}(r_{h})$ is nonmonotonic. It decreases for small $r_{h}$, reaches a single minimum at $r_{h,e_{1}}$, and then increases.
These results confirm our previous analysis that $r_{\rm ps}(r_{h})$ can indeed become nonmonotonic in the isothermal ensemble, which has not been noticed in earlier studies. Interestingly, in the Sch–AdS case ($q=0$), $r_{\rm ps}(r_{h})$ remains monotonic. This can be readily understood from Eq. (41): When $q=0$, the slope $\mathrm{d}P/\mathrm{d}r_{h}$ is not sufficiently negative to drive $\mathrm{d}M/\mathrm{d}r_{h}$ across zero. By contrast, for the RN–AdS case ($q\neq 0$), $P(r_{h})$ exhibits a Van der Waals–like behavior, and at small $r_{h}$ its slope $\mathrm{d}P/\mathrm{d}r_{h}$ becomes strongly negative, which is sufficient to make $\mathrm{d}M/\mathrm{d}r_{h}$ cross zero.
The explicit expression for ${dr_{\mathrm{ps}}}/{dr_{h}}$ reads
@@
For $d=4$ (the RN–AdS case), this reduces to
@@
and, in the Sch–AdS limit ($q=0$), further simplifies to
@@
This coincides with the standard Sch–AdS result, from which ${\mathrm{d}r_{\rm ps}}/{\mathrm{d}r_{h}}>0$ is immediately evident.
To verify the above analysis, we have performed numerical calculations and displayed the results in Fig. 4. In this figure, $T_{c}$ denotes the temperature at the RN–AdS critical point, obtained from the criticality conditions $\partial P/\partial r_{h}=\partial^{2}P/\partial r_{h}^{2}=0$. For the Sch–AdS black hole, $r_{\mathrm{ps}}(r_{h})$ begins at $r_{h,\min}=(d-3)/(4\pi T_{0})$, where $r_{h,\min}$ is defined by $P(r_{h,\min})=0$, and smaller $r_{h}$ would give $P<0$ and are therefore discarded as unphysical. It then increases monotonically, remaining globally monotonic in full agreement with Eq. (51). In contrast, the RN–AdS black hole exhibits the expected non-monotonic behavior: although ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$ at the extremum, the surface gravity remains positive ($\kappa=2\pi T_{0}>0$), indicating that the black hole is nonextremal and the curve need not be truncated.
Figure 4: Behavior of $r_{\rm ps}(r_{h})$ along isothermal paths in $d=4$. The marked point on the curve corresponds to $P=0$ at $r_{h,\min}=(d-3)/(4\pi T_{0})$, and the unphysical portion with $P<0$ is truncated. For the Sch–AdS case, we likewise set the temperature to $T=0.8\,T_{\mathrm{c}}$, where $T_{\mathrm{c}}$ is the critical temperature of the RN–AdS solution with $q=1$. The same choice is adopted for subsequent figures and will not be repeated.
@@ III.2.2 ISCO ($\epsilon=1$)
where $\partial_{P}\Phi_{1}>0$ (see Appendix E for proof). Since van der Waals-type behavior occurs only for charged black holes, we analyze the RN–AdS and Sch–AdS cases separately. Moreover, because the explicit expression for ${\mathrm{d}r_{\mathrm{c}}}/{\mathrm{d}r_{h}}$ is algebraically cumbersome, we focus on whether it changes sign rather than displaying the full formula.
(i) Sch–AdS ($q=0$)
@@
It has a single zero at $r_{h,e_{2}}={(d-3)}/{2\pi T_{0}}$, with ${\mathrm{d}P}/{\mathrm{d}r_{h}}>0$ for $r_{h}<r_{h,e_{2}}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$ for $r_{h}>r_{h,e_{2}}$. Using the criterion (38), together with ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}<0$, $\Gamma_{1}>0$, and $\partial_{P}\Phi_{1}>0$, one thus expects $r_{\rm isco}(r_{h})$ to first decrease and then increase. To make this more explicit, we rewrite the criterion (38) as
@@
For sufficiently large $r_{h}$ one has ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$, which forces $\delta<0$. It remains to verify that $\delta>0$ for small $r_{h}$, ensuring that $\delta$ crosses zero at least once. In this spirit we find
$\begin{cases}\delta\bigl(3r_{h,\min},\,r_{h,\min},\,4,\,0,\,T_{0}\bigr)=\dfrac{1232\,\pi^{2}T_{0}^{2}}{27}>0,&d=4,\\[6.0pt] \delta\bigl(+\infty,\,r_{h,\min},\,d,\,0,\,T_{0}\bigr)=\dfrac{512\,\pi^{3}T_{0}^{3}\,r_{\rm isco}}{(d-3)^{2}(d-1)}>0,&d>4,\end{cases}$ (56)
where $r_{h,\min}=(d-3)/(4\pi T_{0})$ is defined by $P(r_{h,\min})=0$. For $d=4$ and $P=0$, the metric (23) reduces to the Schwarzschild case, giving $r_{\rm isco}=3\,r_{h,\min}$. For $d>4$ and $P=0$, the metric (23) reduces to the higher-dimensional Schwarzschild solution, which admits no ISCO. This corresponds to $\lim_{P\to 0}r_{\rm isco}=+\infty$, see the end of Appendix E for a brief explanation. Equation (56) shows that $\delta>0$ at small $r_{h}$, and together with $\delta<0$ at sufficiently large $r_{h}$ this guarantees that $\delta$ crosses zero, implying that $r_{\rm isco}(r_{h})$ is necessarily nonmonotonic on its full domain.
(ii) RN–AdS ($q\neq 0$)
Since the RN–AdS black hole exhibits a van der Waals–like phase transition, one has
@@
where the equality holds precisely at the critical point $(r_{h,c},T_{c})$. According to the criterion (38), when $r_{h}\geq r_{h,e_{1}}$ (where $r_{h,e_{1}}$ is the zero of $\mathrm{d}M/\mathrm{d}r_{h}$), it follows that $\delta<0$ and hence $\mathrm{d}r_{\mathrm{isco}}/\mathrm{d}r_{h}>0$, so $r_{\mathrm{isco}}(r_{h})$ increases monotonically. For $r_{h}<r_{h,e_{1}}$, however, $\delta$ may become positive. To determine this behavior, we examine the limit $r_{h}\to 0^{+}$.
@@
Meanwhile, from Appendix F, as $r_{h}\to 0^{+}$ one has
@@
This demonstrates that $r_{\mathrm{isco}}(r_{h})$ is nonmonotonic over the full domain of $r_{h}$.
The same conclusion holds for $T<T_{c}$. Along both the small–black hole and large–black hole stable branches, one still has $\mathrm{d}P/\mathrm{d}r_{h}<0$, and the leading-order behavior as $r_{h}\to 0^{+}$ remains identical to the $T\geq T_{c}$ case above. The only difference is that for $T<T_{c}$ an intermediate spinodal (unstable) branch appears, on which $\mathrm{d}P/\mathrm{d}r_{h}>0$. Along this segment $\delta$ could, in principle, change sign, although this appears unlikely and the branch is dynamically skipped by the first-order phase transition. In any case, viewed globally, $r_{\rm isco}(r_{h})$ is necessarily nonmonotonic for every fixed $T_{0}$. These results again confirm our previous analysis that $r_{\rm isco}(r_{h})$ can indeed exhibit nonmonotonic behavior in the isothermal ensemble.
To verify the above analysis, we perform the numerical check, whose outcome is displayed in Fig. 5. It can be seen that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic for both charged and uncharged cases. The distinction lies in that, for the uncharged (Sch–AdS) case, $r_{\mathrm{isco}}(r_{h})$ starts at $r_{h,\min}=(d-3)/(4\pi T_{0})$, whereas for the charged (RN–AdS) black hole, $P(r_{h})>0$ holds throughout the entire domain.
@@
The breakdown of monotonicity naturally raises two questions. First, as noted above, when $r_{c}(r_{h})$ becomes nonmonotonic, can the jump $\Delta_{c}=r_{c}(r_{h,l})-r_{c}(r_{h,s})$ across the phase transition be reduced or even vanish? For RN–AdS black holes, the answer is no. In Appendix G, we prove that the extremal radius $r_{h,e}$ of $r_{\rm ps}(r_{h})$ always lies to the left of the small- and large-black hole radii, $r_{h,s}$ and $r_{h,l}$, namely $r_{h,e}<r_{h,s}<r_{h,l}$. So the phase transition always occurs on the increasing branch, indicating that $r_{\rm ps}$ undergoes a jump together with $r_{h}$ at the phase transition. This behavior is clearly illustrated in Figs. 6 and 7, where the $P(r_{c})$ curve preserves the characteristic S-shaped structure of $P(r_{h})$. But whether this conclusion holds for other black holes must be examined case by case. If instead $r_{h,s}<r_{h,e}<r_{h,l}$, the values of $r_{\rm ps}(r_{h,s})$ and $r_{\rm ps}(r_{h,l})$ could become very close or even identical, potentially weakening or even destroying the reliability of using the circular orbit radius to encode the phase transition. It is therefore essential to analyze, from a theoretical perspective, to what extent phase transitions can be encoded in black hole images. This is precisely the role of the general criterion (21), which allows one to diagnose the monotonicity of $r_{\rm ps}(r_{h})$ in an arbitrary black hole spacetime and thereby assess the reliability of using circular orbit radii to encode phase transitions. Moreover, it is worth noting that, although the proof in Appendix G is presented for the RN–AdS black hole, the underlying method is general and can be applied to other black holes. In particular, we find that the extremal radius of $r_{\rm isco}(r_{h})$ always lies below that of $r_{\rm ps}(r_{h})$. We expect this ordering to persist in other spacetimes.
Second, can nonmonotonic behavior encode additional thermodynamic information in black hole images? Given that both the black hole shadow and accretion-disk images are closely tied to $r_{\rm ps}$, we argue that the answer is positive. A simple idea is that the nonmonotonic evolution of black hole images may distinguish isobaric from isothermal ensembles. Further possibilities will be explored in our future work.
Figure 6: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm ps})$ for $d=4$ and $q=1$. The solid and dashed lines correspond to $P(r_{h})$ and $P(r_{\rm ps})$, respectively. Both sets of lines are plotted for three temperatures: $T=0.8T_{c}$ (red), $T=T_{c}$ (green), and $T=1.2T_{c}$ (blue).
Figure 7: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm isco})$ for $d=4$ and $q=1$. The solid and dashed lines correspond to $P(r_{h})$ and $P(r_{\rm isco})$, respectively. Both sets of lines are plotted for three temperatures: $T=0.8T_{c}$ (red), $T=T_{c}$ (green), and $T=1.2T_{c}$ (blue).